# TEXTOC: Text-driven Object-Centric Style Transfer

### v2

@@ Abstract
We present Text-driven Object-Centric Style Transfer (TEXTOC), a novel method that guides style transfer at an object-centric level using textual inputs. The core of TEXTOC is our Patch-wise Co-Directional (PCD) loss, meticulously designed for precise object-centric transformations that are closely aligned with the input text. This loss combines a patch directional loss for text-guided style direction and a patch distribution consistency loss for even CLIP embedding distribution across object regions. It ensures a seamless and harmonious style transfer across object regions. Key to our method are the Text-Matched Patch Selection (TMPS) and Pre-fixed Region Selection (PRS) modules for identifying object locations via text, eliminating the need for segmentation masks. Lastly, we introduce an Adaptive Background Preservation (ABP) loss to maintain the original style and structural essence of the image’s background. This loss is applied to dynamically identified background areas. Extensive experiments underline the effectiveness of our approach in creating visually coherent and textually aligned style transfers.
Figure 1: Our Object-Centric style transfer results with various image-text scenarios.
* Equal Contribution†Corresponding author
@@ 1 Introduction
In the realm of creative digital industries such as advertising, film, and video game development, the demand for advanced image manipulation is surging. The introduction of object-focused style transfer, driven by textual commands, is transforming these sectors. It allows for detailed and user-friendly adjustments to the visual aspects of objects in images. This innovation empowers designers to bypass traditional manual editing, enabling them to define stylistic alterations using just text, thus facilitating rapid concept development and customization. Consider the ease with which digital fashion elements can be modified, car hues altered for online showrooms, or furniture designs changed in virtual environments, all through simple text instructions, as illustrated in Fig. 1.
@@
Seminal works in style transfer [17, 26, 50, 12, 51, 33, 6] have traditionally relied on reference images to guide the transfer process. While these methods have been foundational, their dependency on visual templates can constrain the creative possibilities. In contrast, a burgeoning wave of research in text-guided style transfer [27, 32, 3, 15, 25, 56] is redefining the landscape. By dispensing with the need for reference images, these novel approaches broaden the horizons for stylistic adaptations, harnessing the power of textual descriptions to steer the creative process. The pioneering approach [3] empowers object-centric editing directly steered by textual descriptions. Yet, this method also presents certain drawbacks; it risks altering the original content’s fidelity, thereby straying from the essential goal of style transfer, which is to enhance without obscuring.
@@
To address this issue, we introduce Text-driven Object-Centric Style Transfer (TEXTOC), a novel approach aimed at transforming the appearance of objects based on textual descriptions. This approach is designed to retain the structural integrity of objects and preserve both the appearance and structure of the background. At the core of TEXTOC lie three pivotal components: the Patch-wise Co-Directional (PCD) loss, the Adaptive Background Preservation (ABP) loss, and the Text-Matched Patch Selection (TMPS) with Pre-fixed Region Selection (PRS) module.
@@
The TMPS and PRS modules are designed to identify and style the locations of objects related to the text. Unlike traditional directional loss [16, 32] that may cause distortion due to a focus on vector directions, our PCD loss, augmented by the TMPS module, directs style transfer in foreground regions aligned with the source text. This ensures a consistent CLIP-embedding distribution across patches from both source and stylized images, promoting a cohesive style transfer within object areas. Our ABP loss is designed to preserve the style and structure of the background in the source image, specifically in non-object areas. It adaptively targets and applies the loss to dynamically detected background regions, ensuring a natural and seamless transition in object-centric style transfer and blending stylized elements with their surroundings as shown in Fig. 1.
@@
  • We present the Text-Matched Patch Selection (TMPS) and Pre-fixed Region Selection (PRS), which identify regions of objects related to text for object-centric style transfer.
  • We propose the Patch-wise Co-Directional (PCD) loss to enable precise style transfer on targeted objects, maintaining the integrity and coherence of the source’s visual aesthetic.
@@ 2.1 Style Transfer
Neural Style Transfer (NST) represents a significant advancement in the domain of image stylization, introduced by [17]. This pioneering approach utilized a pre-trained Convolutional Neural Network (CNN), particularly VGGNet to extract distinct content and style features. However, a notable constraint of NST lies in its substantial computational demand, stemming from the per-image optimization approach. To overcome this, [21] introduce Adaptive Instance Normalization (AdaIN), a technique that aligns the mean and variance of source image features with that of the style image. Building on this, [34, 35] propose Whitening and Coloring Transform, which aligns the entire covariance matrix of the features. This approach results in more refined and superior stylization outcomes. Another significant development is the Unbiased Image Style Transfer [1] that enhances the generalization abilities in style transfer, making it a versatile tool for various style transfer applications.
Figure 2: (Left) Overall pipeline of our TEXTOC consisting of a stylization network (StyleNet), Pre-fixed Region Selection (PRS), Text-Matched Region Selection (TMPS) module and pretrained CLIP encoders. The StyleNet takes a source image $I^{\text{src}}$ and generates an object-wise stylized image $I^{\text{out}}$. The TMPS module is responsible for pinpointing patches that most closely correspond to $T^{\text{src}}$ from the foreground regions identified by the PRS. The selected and augmented patches, $\mathbf{P}^{\text{src}},\mathbf{P}^{\text{out}}$, are then aligned with $T^{\text{src}},T^{\text{tgt}}$ in the CLIP embedding space using Patch-wise Co-Directional (PCD) loss $\mathcal{L}_{\text{pcd}}$. The target text $T^{\text{tgt}}$ is derived by central word selection. Additionally, we apply a content loss $\mathcal{L}_{\text{c}}$, an Adaptive Background Preservation (ABP) loss $\mathcal{L}_{\text{abp}}$ to enhance object-wise style transfer, along with a total variance loss $\mathcal{L}_{\text{tv}}$ for regularization. (Right) Illustration of the functionality of the PCD loss in feature space. It is composed of a patch-wise directional loss $\mathcal{L}_{\text{dir}}$ and a patch distribution consistency loss $\mathcal{L}_{\text{con}}$.
@@ 2.2 Text-Guided Image Synthesis
The Contrastive Language-Image Pretraining (CLIP) model, trained on a large-scale image-text dataset [42], has significantly impacted image synthesis applications [14, 48, 41, 8, 18]. Research [40, 16] building upon the StyleGAN architecture [28, 29] has shown how text descriptors can adapt the style of source images. For instance, the directional loss introduced in [16] showcases remarkable stability and diversity in this domain, overcoming the previous requirement of a style image for transfer. Parallel to these developments, the field has seen a surge in research focusing on diffusion model [19].
@@
2.3 Object-Centric Style Transfer
Recent advancements in object-centric style transfer have seen a surge in customizable techniques. Initial methods [31, 24, 23] rely on reference images for style information, applying style transfer to targeted areas through segmentation networks or additional clustering methods. StyleGAN-NADA [16] marks a paradigm shift in this approach by eliminating the need for reference images and instead utilizing textual directions in the CLIP space for style transfer. This is further advanced by CLIPstyler [32] which successfully applied this novel concept in the style transfer domain. Following this trajectory, the works [54, 27] have introduced style transfer with segmentation map but without the dependency on reference images. A unique approach [3] extracts a relevancy map [5] based on the input text. It allows for stylization focused only on the text-corresponding object regions, moving away from traditional segmentation methods. However, a notable limitation of these techniques is the unintended alteration of both the style and the content information of the object.
@@ 3.1 Overview Framework
The overall pipeline of our model is illustrated in Fig. 2-(Left). We train a stylization network (StyleNet) to generate a stylized image $I^{\text{out}}$, given a source image $I^{\text{src}}$, along with accompanying source text $T^{\text{src}}$, and a style text $T^{\text{sty}}$. We use the text encoder and the image encoder derived from the pre-trained CLIP model, freezing their parameters during the training process. The training of StyleNet incorporates a composite loss function consisting of four distinct components: Patch-wise Co-Directional loss ($\mathcal{L}_{\text{pcd}}$), Adaptive Background Preservation loss ($\mathcal{L}_{\text{abp}}$), a content loss ($\mathcal{L}_{\text{c}}$), and a total variation regularization loss ($\mathcal{L}_{\text{tv}}$). These are weighted by the balance terms $\lambda_{\text{abp}},~{}\lambda_{\text{c}}$ and $\lambda_{\text{tv}}$ as follows:
@@
3.2 Text-Matched Region Selection (TMPS)
A critical aspect of object-centric style transfer is the accurate identification of the object’s location within an image. To address this challenge, we introduce the TMPS module, which is engineered to pinpoint and select image patches correlating to a specified object as described by the source text. Utilizing the image encoder $E_{I}$ and the text encoder $E_{T}$ from the CLIP architecture, TMPS establishes a strong link between the object and its textual descriptor. We define a representative feature vector $f_{avg}$ in Alg. 1, capturing the essential characteristics of the object in the source image. Then, we identify patches $\mathbf{P}^{\text{sel}}$ with features similar to this representative vector. The detailed mechanism of TMPS is explained in Alg. 1.
To streamline the search process for TMPS during the object-centric style transfer, we introduce the Pre-fixed Region Selection (PRS) module within the source image. Applied mainly during the initial iterations, this module delineates a foreground region $M^{\text{fg}}$, configured to coarsely isolate object areas as outlined in Alg. 2. This early demarcation of the object’s location allows for generating patches specifically within the foreground region ($M^{\text{fg}}$) in later iterations. This strategy enhances the precision and efficiency of the style transfer, focusing the modification on the most relevant sections of the image.
@@ 3.2 Text-Matched Region Selection (TMPS)
3.3 Patch-Wise Co-Directional Loss (PCD)
Patch-wise directional loss The foundational concept of directional loss, as initially introduced in [16] and [32], is further refined in our approach. We adapt and advance this concept specifically for object-centric style transfer, with a targeted application on image patches that correspond to the source text $T^{\text{src}}$. We commence by randomly selecting a subset of patches within the foreground region $M^{\text{fg}}$ of the source image. The selected patches, represented as $\mathbf{P}^{\text{src}}\in\{P^{\text{src}}_{1},...,P^{\text{src}}_{N}\}$, are then extracted using our TMPS, as detailed in Alg. 1. Given the patches and the texts, the patch-wise directional loss $\mathcal{L}_{\text{dir}}$ is defined as follows:
@@ 3.3 Patch-Wise Co-Directional Loss (PCD)
Target text generation To formulate the target text $T^{\text{tgt}}$, we apply central word selection technique $\text{Cen}(\cdot)$ in (3) leveraging Spacy [20], and merge the source text $T^{\text{src}}$ and the style text $T^{\text{sty}}$. This design is rooted in the manifold augmentation technique proposed in [55] leveraging text embeddings from models like BERT [9], GPT [43], and CLIP. This amalgamation of texts aligns more closely with the user’s intended styling objectives. For instance, to adapt the style of ‘red apple’ to appear green, we engage TMPS using ‘red apple’ but modify the target text to ‘green apple’, with an emphasis on ‘apple’ as the central word. This refined approach to text input in TMPS enables more accurate patch selection, facilitating precise object-centric style transfer.
Patch distribution consistency loss Traditional directional loss functions typically emphasize the direction of vectors while often neglecting their semantic information. This emphasis can inadvertently lead to a misalignment between patches in the source image and those in the stylized image as depicted in Fig. 2-(Right). Such a discrepancy can lead to the collapse of semantic information, resulting in a distorted transformation that causes the loss of information from the source image, contradicting the fundamental aim of style transfer. To address this issue, we design the patch distribution consistency loss $\mathcal{L}_{\text{con}}$ as follows:
@@
ZeCon [53]      0.23 7.49 0.28 5.33 0.63 0.64 0.24 19.35
@@
Table 1: Quantitative comparison of our method with the recent text-guided style transfer methods using foreground quality metrics and background quality metrics. The symbols $\uparrow$ and $\downarrow$ indicate higher values are better and lower values are better, respectively.
@@ 3.4 Adaptive Background Preservation (ABP) Loss
We introduce the ABP loss, designed to ensure the non-stylization of the background region. Our approach begins by identifying a pre-fixed foreground $M^{\text{fg}}$, as outlined in Alg. 2. In each iteration, the adaptive foreground mask $*M^{\text{fg*}}$ and background mask $*M^{\text{bg*}}$ are dynamically updated as follows:
@@
This loss enables a well-balanced integration of the stylized and original regions within the image. Specifically, while the foreground areas undergo stylization, the background’s integrity is preserved.
Figure 3: Comparison of our method with various text-guided style transfer models. To the left of the solid line are the qualitative results of our model and non-diffusion based models, to the right are the results from diffusion-based methods.

### v3

# Style-Editor: Text-driven object-centric style editing

@@ Abstract
We present Text-driven object-centric style editing model named Style-Editor, a novel method that guides style editing at an object-centric level using textual inputs. The core of Style-Editor is our Patch-wise Co-Directional (PCD) loss, meticulously designed for precise object-centric editing that are closely aligned with the input text. This loss combines a patch directional loss for text-guided style direction and a patch distribution consistency loss for even CLIP embedding distribution across object regions. It ensures a seamless and harmonious style editing across object regions. Key to our method are the Text-Matched Patch Selection (TMPS) and Pre-fixed Region Selection (PRS) modules for identifying object locations via text, eliminating the need for segmentation masks. Lastly, we introduce an Adaptive Background Preservation (ABP) loss to maintain the original style and structural essence of the image’s background. This loss is applied to dynamically identified background areas. Extensive experiments underline the effectiveness of our approach in creating visually coherent and textually aligned style editing.
Figure 1: Results of our Style-Editor under diverse textual conditions. To the left and right of the arrow($\rightarrow$) indicate the source text $T^{\text{src}}$ and style text $T^{\text{sty}}$, respectively.
11footnotetext: Equal contribution.22footnotetext: Corresponding author.
@@ 1 Introduction
In the realm of creative digital industries such as advertising, film, and video game development, the demand for advanced image manipulation is surging. The introduction of an object-focused style editing model, driven by textual commands, is transforming these sectors. It allows for detailed and user-friendly adjustments to the visual aspects of objects in images. This innovation empowers designers to bypass traditional manual editing, enabling them to define stylistic alterations using just text, thus facilitating rapid concept development and customization. Consider the ease with which digital fashion elements can be modified, car hues altered for online showrooms, or furniture designs changed in virtual environments, all through simple text instructions, as illustrated in Fig. 2.
Figure 2: Our editing results in industrial applications.
@@
Style editing began with the foundational concept of style transfer, which traditionally relies on reference images to guide the transformation process, as demonstrated by seminal works [17, 28, 57, 12, 58, 35, 6]. While these methods have been foundational, their dependency on visual templates can constrain creative possibilities. In contrast, a burgeoning wave of research in text-guided style transfer [29, 34, 2, 15, 27, 66] is redefining the landscape. By dispensing with the need for reference images, these novel approaches broaden the horizons for stylistic editing, harnessing the power of textual descriptions to steer the creative process. The pioneering approach [2, 4, 43, 56, 40] empowers object-centric editing directly steered by textual descriptions. Yet, this method also presents certain drawbacks; it risks altering the original content’s fidelity, may fail to accurately capture the intended textual descriptions, or may unintentionally edit undesired parts of the image.
@@
To address this issue, we introduce Text-driven object-centric style editing model named Style-Editor, a novel approach aimed at transforming the appearance of objects based on textual descriptions. This approach is designed to retain the structural integrity of objects and preserve both the appearance and structure of the background. At the core of Style-Editor lie three pivotal components: the Patch-wise Co-Directional (PCD) loss, the Adaptive Background Preservation (ABP) loss, and the Text-Matched Patch Selection (TMPS) with Pre-fixed Region Selection (PRS) module.
@@
Leveraging the robust zero-shot image classification capabilities of the CLIP model, the TMPS and PRS modules are designed to identify and style the locations of objects related to the text. Unlike traditional directional loss [16, 34, 63] that may cause distortion due to a focus on vector directions, our PCD loss, augmented by the TMPS module, directs style editing in foreground regions aligned with the source text. This ensures a consistent CLIP-embedding distribution across patches from both source and stylized images, promoting a cohesive style editing within object areas. Our ABP loss is designed to preserve the style and structure of the background in the source image, specifically in non-object areas. It adaptively targets and applies the loss to dynamically detected background regions, ensuring a natural and seamless transition in object-centric style editing and blending stylized elements with their surroundings as shown in Fig. 2.
@@
  • We present the Text-Matched Patch Selection (TMPS) and Pre-fixed Region Selection (PRS), which identify regions of objects related to text for object-centric style editing.
  • We propose the Patch-wise Co-Directional (PCD) loss to enable precise style editing on targeted objects, maintaining the integrity and coherence of the source’s visual aesthetic.
@@ 2.1 Style Transfer
Neural Style Transfer (NST) represents a significant advancement in the domain of image stylization, introduced by [17]. This pioneering approach utilized a pre-trained Convolutional Neural Network (CNN), particularly VGGNet, to extract distinct content and style features. However, a notable constraint of NST lies in its substantial computational demand, stemming from the per-image optimization approach. To overcome this, [22] introduced Adaptive Instance Normalization (AdaIN), a technique that aligns the mean and variance of source image features with those of the style image. Building on this, [36, 37] proposed the Whitening and Coloring Transform (WCT), which aligns the entire covariance matrix of the features, resulting in more refined and superior stylization outcomes. With the advent of attention mechanisms in neural networks [59, 11], new style transfer models have emerged that utilize these mechanisms to achieve impressive results [39, 20, 8, 42, 64], reflecting the ongoing evolution of style transfer technology.
Figure 3: (Left) Overall pipeline of our Style-Editor consisting of a style editing network (StyleNet), Pre-fixed Region Selection (PRS), Text-Matched Patch Selection (TMPS) module and pretrained CLIP encoders. The StyleNet takes a source image $I^{\text{src}}$ and generates an object-wise stylized image $I^{\text{out}}$. The TMPS module is responsible for pinpointing patches that most closely correspond to $T^{\text{src}}$ from the foreground regions identified by the PRS. The selected and augmented patches, $\mathbf{P}^{\text{src}},\mathbf{P}^{\text{out}}$, are then aligned with $T^{\text{src}},T^{\text{tgt}}$ in the CLIP embedding space using Patch-wise Co-Directional (PCD) loss $\mathcal{L}_{\text{pcd}}$. The target text $T^{\text{tgt}}$ is derived by central word selection. Additionally, we apply a content loss $\mathcal{L}_{\text{c}}$, an Adaptive Background Preservation (ABP) loss $\mathcal{L}_{\text{abp}}$ to enhance object-centric style editing, along with a total variance loss $\mathcal{L}_{\text{tv}}$ for regularization. (Right) Illustration of the functionality of the PCD loss $\mathcal{L}_{\text{pcd}}$ in feature space. It is composed of a patch-wise directional loss $\mathcal{L}_{\text{dir}}$ and a patch distribution consistency loss $\mathcal{L}_{\text{con}}$.
@@ 2.2 Text-Guided Image Synthesis
The Contrastive Language-Image Pretraining (CLIP) model, trained on a large-scale image-text dataset [49], has significantly impacted image synthesis applications [47, 55, 18, 14]. Research [45, 16] building upon the StyleGAN architecture [30, 31] has shown how text descriptors can adapt the style of source images. For instance, the directional loss introduced in [16] showcases remarkable stability and diversity in this domain, overcoming the previous requirement of a style image for transfer. Parallel to these developments, the field has seen a surge in research focusing on diffusion model [19].
@@
2.3 Text-driven Object-Centric Style Editing
Style editing modifies the color or texture of the original image to match a desired style while preserving the image’s content [24]. Early methods [33, 26, 25] utilized reference images for style information, applying style editing to targeted areas through segmentation networks or clustering methods. StyleGAN-NADA [16] introduced a paradigm shift by eliminating the need for reference images and instead leveraging textual directions in the CLIP space for domain transfer. Building on this, CLIPstyler [34] successfully applied this concept in text-driven style editing, advancing this field. Following this trajectory, [2] proposed a text-driven object-centric style editing approach by extracting a relevancy map [5].
Another approach to text-driven object-centric style editing involves diffusion models trained on large vision-text datasets, such as [50, 54]. Within these approaches, some involve training the model on newly generated text-image pairs [4] or optimizing null-text [40]. Additionally, methods [56, 43, 3, 46] have been developed to edit feature maps of the diffusion model, enabling more precise control over style editing based on specific textual guidance. Despite these diverse advancements, a notable limitation remains: these techniques often lead to unintended alterations of both style and content, and they often fail to achieve high levels of style editing fidelity.
@@ 3.1 Overview Framework
The overall pipeline of our model is illustrated in Fig. 3-(Left). We train a style editing network (StyleNet) to generate a stylized image $I^{\text{out}}$, given a source image $I^{\text{src}}$, along with accompanying source text $T^{\text{src}}$, and a style text $T^{\text{sty}}$. We use the text encoder and the image encoder derived from the pre-trained CLIP model, freezing their parameters during the training process. The training of StyleNet incorporates a composite loss function consisting of four distinct components: Patch-wise Co-Directional loss ($\mathcal{L}_{\text{pcd}}$), Adaptive Background Preservation loss ($\mathcal{L}_{\text{abp}}$), a content loss ($\mathcal{L}_{\text{c}}$), and a total variation regularization loss ($\mathcal{L}_{\text{tv}}$). These are weighted by the balance terms $\lambda_{\text{abp}},~{}\lambda_{\text{c}}$ and $\lambda_{\text{tv}}$ as follows:
@@
3.2 Text-Matched Patch Selection (TMPS)
A critical aspect of object-centric style editing is the accurate identification of the object’s location within an image. To address this challenge, we introduce the TMPS module. This module pinpoints and selects image patches that correlate to a specified object as described by the source text, leveraging the zero-shot image classification capability of the CLIP model. Utilizing the image encoder $E_{I}$ and the text encoder $E_{T}$ from the CLIP architecture, TMPS establishes a strong link between the object and its textual descriptor. We define a representative feature vector $f_{\text{avg}}$ in Alg. 1, capturing the essential characteristics of the object in the source image. Then, we identify patches $\mathbf{P}^{\text{sel}}$ with features similar to this representative vector. The detailed mechanism of TMPS is explained in Alg. 1.
To streamline the search process for TMPS during the object-centric style editing, we introduce the Pre-fixed Region Selection (PRS) module within the source image. Applied during the initial iterations, this module delineates a foreground region $M^{\text{fg}}$, configured to coarsely isolate object areas as outlined in Alg. 2. This early demarcation of the object’s location allows for generating patches specifically within the foreground region ($M^{\text{fg}}$) in later iterations. This strategy enhances the precision and efficiency of the style editing, focusing the modification on the most relevant sections of the image.
@@ 3.2 Text-Matched Patch Selection (TMPS)
3.3 Patch-Wise Co-Directional Loss (PCD)
Patch-wise directional loss The foundational concept of directional loss, as initially introduced in [16] and [34], is further refined in our approach. We adapt and advance this concept specifically for object-centric style editing, with a targeted application on image patches that correspond to the source text $T^{\text{src}}$. We commence by randomly selecting a subset of patches within the foreground region $M^{\text{fg}}$ of the source image. The selected patches, represented as $\mathbf{P}^{\text{src}}\in\{P^{\text{src}}_{1},...,P^{\text{src}}_{N}\}$, are then extracted using our TMPS, as detailed in Alg. 1. Given the patches and the texts, the patch-wise directional loss $\mathcal{L}_{\text{dir}}$ is defined as follows:
Target text generation To formulate the target text $T^{\text{tgt}}$, we apply central word selection technique $\text{Cen}(\cdot)$ in (3) leveraging Spacy [21], and merge the source text $T^{\text{src}}$ and the style text $T^{\text{sty}}$. This design is rooted in the manifold augmentation technique proposed in [65] leveraging text embeddings from models like BERT [9], GPT [48], and CLIP. This amalgamation of texts aligns more closely with the user’s intended styling objectives. For instance, to adapt the style of ‘red apple’ to appear green, we engage TMPS using ‘red apple’ but modify the target text to ‘green apple’, with an emphasis on ‘apple’ as the central word. This refined approach to text input in TMPS enables more accurate patch selection, facilitating precise object-centric style editing.
Patch distribution consistency loss Traditional directional loss functions typically emphasize the direction of vectors while often neglecting their semantic information. This emphasis can inadvertently lead to a misalignment between patches in the source image and those in the stylized image as depicted in Fig. 3-(Right). Such a discrepancy can lead to the collapse of semantic information, resulting in a distorted transformation that causes the loss of information from the source image, contradicting the fundamental aim of style editing. To address this issue, we design the patch distribution consistency loss $\mathcal{L}_{\text{con}}$ as follows:
@@ 3.3 Patch-Wise Co-Directional Loss (PCD)
pix2pix-zero [43]      0.22 8.12 0.40 4.91 1.07 0.61 0.26 17.11
LEDITS++ [3]      0.22 6.81 0.18 2.92 0.44 0.74 0.14 21.66
local-prompt-mixing [46]      0.20 9.65 0.24 4.65 0.61 0.67 0.18 20.01
StylerDALLE [62]      0.26 8.35 0.51 5.71 1.22 0.52 0.28 15.42
@@
Table 1: Quantitative comparison of our method with the recent text-guided style editing methods using foreground quality metrics and background quality metrics. The symbols $\uparrow$ and $\downarrow$ indicate higher values are better and lower values are better, respectively.
@@ 3.4 Adaptive Background Preservation (ABP) Loss
We introduce the ABP loss, designed to ensure the non-style editing of the background region. Our approach begins by identifying foreground regions as outlined in Alg. 1, Alg. 2. In each iteration, the adaptive foreground mask $*M^{\text{fg*}}$ and background mask $*M^{\text{bg*}}$ are dynamically updated as follows:
@@
This loss enables a well-balanced integration of the stylized and original regions within the image. Specifically, while the foreground areas undergo style editing, the background’s integrity is preserved.
Figure 4: Comparison of our method with various text-guided style editing models. To the left of the solid line are the qualitative results of our model and non-diffusion based models, to the right are the results from diffusion-based methods.